### v1

# Reffly: Melody-Constrained Lyrics Editing Model

Abstract
Automatic melody-to-lyric generation aims to produce lyrics that align with a given melody. Although previous work can generate lyrics based on high-level control signals, such as keywords or genre, they often struggle with three challenges: (1) lack of controllability, as prior works are only able to produce lyrics from scratch, with little or no control over the content; (2) inability to generate fully structured songs with the desired format; and (3) failure to align prominent words in the lyrics with prominent notes in the melody, resulting in poor lyrics-melody alignment. In this work, we introduce Reffly (REvision Framework For Lyrics), the first revision framework designed to edit arbitrary forms of plain text draft into high-quality, full-fledged song lyrics. Our approach ensures that the generated lyrics retain the original meaning of the draft, align with the melody, and adhere to the desired song structures. We demonstrate that Reffly performs well in diverse task settings, such as lyrics revision and song translation. Experimental results show that our model outperforms strong baselines, such as Lyra Tian et al. (2023) and GPT-4, by 25% in both musicality and text quality.
**footnotetext: Equal contribution
1 Introduction
Figure 1: Human singers naturally emphasize certain words when singing, which align with prominent notes to ensure musical flow (details in §$\S$ 3.1). However, LLMs like GPT-4 often misalign these prominent words (e.g."arm" with non-prominent notes) or omit syllables (e.g. no words for the last four notes), lowering lyric quality(§$\S$ 2). Our model, Reffly, refines less singable drafts into melodically-aligned lyrics while preserving the meaning. Listen to the audios for an intuitive sense.
Automatic melody-to-lyric (M2L) generation is an emerging task focuses on creating lyrics that align with the rhythm and structure of a given melody (Sheng et al., 2021). Existing M2L works usually produce lyrics from scratch with little user input Sheng et al. (2021); Tian et al. (2023); Ding et al. (2024), limiting the incorporation of user communication intent. While there are many practical applications revising plain texts to lyrics, under the context of lyrics translation Guo et al. (2022), style transfer (e.g., generating rap lyrics from news articles Nikolov et al. (2020)), and AI-assisted lyric refinement Zhang et al. (2022b). They did not consider melody constraints and thus cannot generate lyrics that are singable with a given melody. A revision framework which edits plain text into melody-aligned lyrics is yet to be explored.
Furthermore, state-of-the-art LLMs et al. (2023) struggle with producing singable lyrics that align precisely with a specific melody. For instance, as illustrated in Figure 1, GPT-4 generates coherent lyrics but fails to synchronize syllables with the last four music notes. Moreover, important words like ‘arm’ and ‘eye’ are paired with less significant notes, disrupting the overall musical flow and resulting in low prosody. On the other hand, prior works on M2L generation Zhang et al. (2022a); Ram et al. (2021) either perform subpar on lyric quality, or/and struggle to generate full-structured lyrics, such as verses or choruses, with good singability. For example, Sheng et al. (2021) employed masked sequence-to-sequence pre-training and attention-based alignment modeling but overlooked prosody and singability, which limited lyrical quality. While recent research (Tian et al., 2023) has enhanced singability by aligning stressed syllables with prolonged musical notes, it still fails to consider prosody, which emphasizes the alignment of key lyrical words and prominent musical notes – a crucial aspect for lyrics-melody alignment.
Addressing these challenges, we propose a novel revision framework, Reffly, which transfers a draft prose to structured and singable lyrics align with a piece of melody. To enhance melody-lyric alignment, we develop a training-free heuristic for capturing prominent lyrical words and musical notes (§$\S$3.1). Since the melody-aligned lyric data is scarce due to copyright constraints, we design an instruction-based mechanism to guide LLMs towards highly singable lyrics by training on a synthetic dataset (§$\S$ 3). Reffly can generate full-length songs with lyrical refrains of verse (develops the song’s plot and message) and chorus (the repeated, memorable musical motif).
Our contributions are summarized as follows:
  • We propose the first melody-constrained lyric revision framework that, given a predefined melody, transfers an arbitrary text (also referred to as a draft or unsingable lyrics) to a full-length, melody-aligned lyrics with high singability and prosody (also referred to as revised or singable lyrics).
  • We introduce a training-free heuristic for capturing melody-lyrics alignment, semantically and musically, to improve both singability and prosody. Correspondingly, we also contribute a expert labeled dataset with fine-grained annotations of music sheets.  [^1] Anonymous dataset source: https://bit.ly/3X6nCquhttps://bit.ly/3RytMfR
  • In comprehensive experiments across two settings: 1) generation of lyrics from user-specified inputs, and 2) translation of lyrics from Chinese to English, Reffly significantly enhances lyrics-melody alignment and text quality of the generated lyrics, resulting in a 25% and 34% improvements over strong baselines in terms of musicality and overall preference, respectively. [^2] Anonymous demo: https://bit.ly/4fGKWT3
2 Problem Setup and Background
2.1 What Makes a Good Lyric?
Great lyrics harmonize with the melody, blending musicality (e.g., singability, prosody) with textual quality (e.g., coherence, creativeness) Perricone (2018). Here, we elaborate the two terms related to musicality below:
  • Singability is what makes a song easier to sing. For example, it is considered not singable when one single music note maps to a multi-syllable word (e.g., beau-ti-ful) in the lyrics.
  • Prosody measures whether melody and lyrics rise and/or fall together Perricone (2018). Lyrics with good prosody highlights prominent words by matching them with prominent notes. For example, in Figure 1, Reffly enhances expression by stressing prominent words like ‘embrace’ and ‘eye’ by aligning them with prominent notes.
These concepts guided the development of heuristics to better align lyrics with the melody.(§$\S$ 3.1).
2.2 Task Formulation
Goal
Given a predefined melody and a plain-text draft, our goal is to revise the unsingable draft into full-length lyrics that excel in both musicality and textual quality.
Formulation
[…]
Figure 2: The overview of the inference process of Reffly, an iterative approach to revise each sentence from the unsingable draft based on corresponding music constraint that is extracted from the music score. Reffly begins by taking the melody and the unsingable draft as inputs. It then extracts features and constructs a prompt (Steps 1 and 2). Subsequently, it prompts a trained revision module to revise the unsingable draft (Step 3.1) and aligns the revised draft with the melody constraints using an alignment algorithm (Step 3.2). Note that only Step 3.1 requires training, and all other processes are training-free.
3 Revision Framework for Lyrics
Figure 2 illustrates the inference process of Reffly. To manage the complexity of lyric revision, the revision process is conducted at the sentence level. We iteratively revise each sentence from the unsignable draft to lyric that fits the melody, aligning the prominent words with prominent notes while maintaining the overall coherence.
In this section, we detail each component of Reffly. We developed a training-free heuristic for capturing prominent lyrical words and musical notes (Section §$\S$3.1). Then, §$\S$3.2 introduces revision module for refining unsingable drafts based on musical constraints. Next, §$\S$3.3 provides an overview of the inference process to achieve optimal lyric-melody alignment. Lastly, §$\S$3.4 describes our methods for controlling song structure and generating full-length songs.
3.1 Prominent Musical Notes and Words
Experienced human singers naturally stress prominent words for expressiveness, which align with prominent musical notes Robinson (2005). We introduce the heuristic for identifying prominent notes and words, which are used in generating singable lyric (§$\S$3.3) and the construction of synthetic training data (§$\S$4.1.1).
Prominent Musical Notes
Inspired by prior research in music theory Caroline Palmer (2006), we developed a heuristic to identify prominent musical notes that stand out in melodies. Our heuristic identifies prominent notes based on three fundamental characteristics of music: Time signature, Rhythm, and Pitch. See Appendix E.2 for details.
Prominent words
Nouns, verbs, adjectives are crucial for effectively conveying meaning through emphasis and recognition (Reikofski, 2015). We validated this concept within the context of lyrics using our constructed datasets, which reveal that nouns, verbs, and adjectives are the most significant parts of speech in terms of prominence: 91.8% of important words are either nouns, verbs, or adjectives. Hence, we identify these part-of-speeches as prominent words (details in Appendix E.3).
3.2 Revision Module
To achieve high-quality lyrics revision, we fine-tuned a revision module with an instruction-based mechanism, as shown in Figure 2, We selected LlaMA2-13b-chat Touvron et al. (2023) as our base model. To effectively transform an unsingable draft into lyrics that fit a given melody while preserving the original meaning, we must address three main tasks:
(1) Syllable planning, which involves generating sentences with the required number of syllables for singability. This task is challenging for LLMs due to difficulties with numerical planning Sun et al. (2023) and the lack of a specific fine-tuning dataset. Moreover, aligning prominent words with prominent notes requires defining which notes and words are prominent.
(2) Aligning prominent words with prominent notes to enhance prosody by ensuring that all prominent words match prominent notes. We propose a "pseudo music constraint" and an instruction template (as illustrate in Figure 3) to improve the syllable planning and word-note matching. The pseudo constraint, derived from generated lyrics, indicates prominent note positions and syllable counts. During training, the model follows pseudo constraints, while in inference, it applies melody constraints. (refer §$\S$§D.2 for more details)
(3) Maintaining coherence for fluent, logical lyrics. Our revision approach operates at the sentence level. To ensure coherence, we provide the model with previously generated lyrics and the song’s title for each training data point.
3.3 Generate Singable Lyric
To generate a full-length singable lyric, we use Reffly to iteratively revise each unsingable sentence. These revised sentences are then seamlessly merged to form the final output.
As illustrated in figure 2, Reffly takes the melody and each sentence from unsingable draft as input, to produce singable lyrics as the output, with three steps as following.
Step 1: Extract Features.
We extract features from input for further processing. We use our proposed heuristic (refer to §$\S$3.1) to identify prominent musical notes from the melody, encoding them into a melody constraint. We then assemble a prompt using the constraint, title, context, and lyrics draft as input features.
Step 2: Assemble Prompt.
The prompt includes instructions regarding the syllable count to match the melody’s rhythm, steps for refining lyrics, consistency checks with previous context to ensure coherence, and song structure guidance to maintain desired song formats.
Step 3: Revise and Align.
To enhance the model capability of generating singable and prominence-aligned lyrics, we break down the process into two sub-steps, and iteratively generate the lyric.
Step 3.1 Revise: We use diverse beam search to generate multiple candidate revisions of the unsingable draft, evaluating each for singability. A lyric is singable if: 1) each note corresponds to one or zero syllables; 2) each syllable in multi-syllable words matches a note no longer than a half-note; 3) multi-syllable words do not cross rests. If no candidates meet these criteria, we restart the process with a rephrased draft.
1:input: List of candidates $C$, orignal draft $o$, melody constraint $m$, max number of ties $K$
2:output: Revised singable lyric
3:
4:$c_{qualified}$ = empty list
5:for candidate $c$ in $C$ do
6:     $c_{num}=$ calculate_num_ties$(c,m)$
7:     if  $0\leq c_{num}\leq K$ then
8:         $c_{tie}=add\_tie(c,m,c_{num})$
9:         $c_{qualified}+=c_{tie}$
10:     end if
11:end for
12:for candidate $c$ in $c_{qualified}$ do
[…]
14:end for
15:return $c_{best}$
Algorithm 1 Candidates Selection
Step 3.2 Align Unsingable Draft: Algorithm 1 illustrates the alignment algorithm. For each qualified candidate $c$ and music constraint $m$, we determine the number of ties (a common musical notation that maps more than one notes to one syllable) to add using the following:
#$\vspace{-2mm}\#\text{Ties}=\#\text{Notes}(m)-\#\text{Syllables}(c)\vspace{-2mm}$ (1)
Next, we define $K$, a tune-able hyper parameter, as maximum number of ties allowed within each musical phrase. We set $K=2$ as a reasonable number in all of our experiments. If #$\#\text{ties}<0$ or #$\#\text{ties}>K$, we reject the input. Otherwise, we explore all feasible positions to insert ties, aiming to maximize the number of prominent words mapped to prominent notes.
Finally, we select the candidate whose most important words align with prominent notes. If multiple candidates align perfectly, we choose the one most similar to the original sentence based on BERTScore. Zhang et al. (2019).
3.4 Incorporating Full-Song Structure
To generate full-length songs, our framework integrates structure-awareness. During the fine-tuning phase, we get song structure (as introduced in section 2.1) for each lyric line from data source, enabling the model to identify distinctive features of song parts like verses and choruses. As shown in Figure 3, structure tags are embedded directly within the instructions component of our training sequences. Then, as illustrated in Figure 2, during inference (Step 2) , the target structure tags are integrated into the prompt. This strategic incorporation ensures that the generated lyrics align with the melody and are cohesively structured.
Figure 3: An exemplary data point in the fine tune dataset. The task is to use a rephrased input, title, music constraint, previously generated lyrics, and assambled instruction to generate the original lyrics, and some explanation. Rephrasing is done by GPT-3.5. During training, the revision model is guided by pseudo melody constraints derived from the original lyrics, enabling it to follow real melody constraints during inference.
4 Experiments setup
4.1 Dataset Setup
We prepared two datasets: 1) a synthetic training dataset, constructed using our devised strategy to fine-tune the model, and 2) a validation dataset, employed to assess our heuristic (§$\S$ 3.1).
4.1.1 Training dataset
As shown in Figure 3, the objective of our training dataset is to instruct the model to generate original lyrics by revising draft lyrics, following music constraints. We construct this dataset using 3,500 song-lyrics collected from the internet. Notably, our revision model only requires lyrics during training, alleviating the lack of aligned melody-lyrics data and potential copy-right issue. [^4] More details about the different components of input and output of training dataset can be found in Appendix D.2.
4.1.2 Validation dataset
To evaluate the accuracy of our prominent note extraction heuristic, we created a dataset consisting of 100 song clips annotated by professional musicians, highlighting all the important notes in each melody. This dataset encompasses a diverse range of music styles, including Jazz, Country, Blues, Folk, Pop, and Comedy. An exemplary data point is shown in Figure 6 at appendix.
4.2 Tasks Setup
Our model’s versatility is demonstrated through its performance across three distinct tasks. We prompt LLaMA2-13b in a few-shot manner to generate lyrics drafts based on user thoughts [^5] Details on generating the lyrics draft are in Appendix C.2:
1. Lyrics generation from arbitrary content. This task generates song lyrics from scratch, starting with a draft based on scattered user thoughts. The lyrics’ quality and melody alignment are evaluated using automated metrics and human judgment.
2. Full-Length generation with song structures (Structure-Aware Generation) In this task, we generate lyrics with specific structural requirements, starting from scattered user feedback. Domain experts then review these generated lyrics for coherence and clarity.
3. Song Translation: This task focuses on translating lyrics from Chinese to English. The initial draft is a straightforward text translation produced by a translation model. We recruit bilingual evaluators to assess the translated lyrics quality.
4.3 Baselines
We compare our frameworks performance with following models: 1. Lyra is an unsupervised, hierarchical melody-conditioned lyric generator that can generate high-quality lyrics with content control without training on melody-lyric data Tian et al. (2023). 2. SongMass is an LLM design that leveraging masked sequence to sequence (MASS) pre-training and attention based alignment modeling for lyric-to-melody and melody-to-lyric generation Sheng et al. (2021). 3. GPT-4 is a strong versatile LLM et al. (2023) to compare with. We utilize few-shot prompt to provide a template and instruct the model to follow it. 4. Reffly-S represents a variant of our proposed model, from which we have removed the candidate selection algorithm ( introduced in section 3 ). 5. Reffly-I is a variation of our proposed model, where we exclude the instruction component during training(§$\S$ 4.1.1 ). [^6] We describe details of baselines in Appendix C.1.
4.4 Automatic Evaluation Setup
We evaluate the generated lyrics on text quality and melody alignment. For text quality, we assess several aspects: 1) Diversity, measured by calculating the number of unique n-grams in the text; 2) Perplexity, using GPT-2 to evaluate fluency and predictability; 3) Similarity, evaluated with BERTScore, to measure the similarity between our model-generated lyrics and the lyrics draft. For melody alignment, we proposed the prominent word-note matching rate, as explained in §$\S$ 2.1, to measure how well prominent words are aligned with prominent musical notes.
4.5 Human Evaluation Setup
Qualification Task is conducted to select qualified annotators with 1) adequate expertise in song and lyric annotation, and 2) reliable focus on the Mechanical Turk platform. Additional details on the qualification task are provided in Appendix B.1.
Annotation Task Our annotation process is comparative, with annotators reviewing groups of songs produced by various systems that share the same melody and title. All baseline models were assessed. At least three workers annotated each song, rating the lyrics’ quality on a 1-5 Likert scale across five categories. For musicality, the workers assessed prosody (whether prominent words were exaggerated by melody), intelligibility (whether the lyric content was easy to understand when listen to it), and singability (how clearly the lyrics could be understood). In terms of text quality, they evaluated coherence and creativity. The average inter-annotator agreement in terms of Pearson correlation was 0.69.
Automatic Evaluation Human Evaluation
Model
Diversity
(Unigram)$\uparrow$
Diversity
(Bigram)$\uparrow$
Similar-
$\uparrow$
Perplex-
$\downarrow$
Match
[…]
Intelli-
$\uparrow$
Singab-
$\uparrow$
Creati-
$\uparrow$
Lyra 0.52 0.86 0.72 3305 0.48 1.97 1.66 2.02 1.83 1.70
SongMASS 0.50 0.76 3759 0.40 1.35 1.11 1.65 1.46 1.07
[…]
Table 1: Evaluation Results for the Arbitrary Generation task. Reffly and its variants (Reffly-S and Reffly-I) consistently outperform other models across most metrics, both in automatic and human evaluations.
5 Results
Note Extraction
Success Rate
Alignment
Success Rate
Duration-Only 74% 43%
[…]
Comprehensive 96% 91%
Table 2: Comparison of three extraction and alignment strategies. The highest performance in each category is highlighted in bold, illustrating the superior effectiveness of our strategy in both note extraction (96%) and alignment (91%).
5.1 Effectiveness of the heuristic
Our heuristic for identifying prominent notes and words is validated against baselines using a musician-annotated dataset of 100 song clips (§$\S$ further details in 4.1.2). The first baseline relies solely on note duration to determine prominence, similar to the decoding constraints used in Lyra Tian et al. (2023), and pairs this with our word extraction heuristic. The second baseline utilizes our note extraction heuristic but restricts prominent words to nouns and verbs.
As shown in Table 2, our prominent note extraction heuristic achieves an accuracy of 96%, substantially outperforming both baselines. Furthermore, our comprehensive heuristic for extracting prominent words and notes yields a 91% successful alignment rate  [^7] Successful alignment rate is calculated as the proportion of prominent words correctly mapped to prominent notes., surpassing the best baseline by 26%. These results underscore the effectiveness and non-trivial nature of our approach in capturing the alignment between prominent words and prominent notes (more details in Appendix E.4).
5.2 Result of Lyrics Generation from Arbitrary Content
The results of automatic evaluation (mainly assesses fluency, topic relevance, and melody-lyric alignment) and human evaluations (assesses overall quality across multiple aspects such as musicality, creativity, etc.) are reported in Table 1.
Automatic Results
The similarity scores in Table 1 indicate that Reffly and GPT-4 excel in preserving the meaning of unsingable drafts. In contrast, SongMASS and Lyra surpass GPT-4 in terms of musicality, but at the cost of fluency. The qualitative example (shown in Section 6) shows that SongMASS and Lyra tend to generate cropped sentences to fit the music, leading to higher perplexity. Although GPT-4 matches Reffly in retaining lyrical meaning, it falls short in diversity and melody alignment, as reflected in its lowest Match Rate. Overall, Reffly surpasses all baselines, producing lyrics with superior textual quality, optimal melody alignment, and faithful preservation of the original draft’s meaning.
Automatic Evaluation Human Evaluation
Model
[…]
Perplex-
$\downarrow$
Match
[…]
Coher-
$\uparrow$
Intelli-
$\uparrow$
Singab-
$\uparrow$
Creati-
$\uparrow$
Translate
[…]
GPT-4 0.50 0.76 522 0.35 1.59 2.24 1.83 1.54 2.26 2.33
Reffly 0.59 0.87 310 0.83 3.28 3.08 3.25 3.08 2.69 3.04
Table 3: Song translation task result. Reffly scores the highest for all metrics.
GPT-4 Reffly
Prosody 1.31 3.10
Sinability 1.28 3.11
Coherence 1.84 2.70
Creativity 1.85 2.62
Intelligibility 1.26 2.99
Structural Clarity 1.15 3.36
Table 4: Structure-aware generation results: Reffly outperforms GPT-4 by producing lyrics with a clearer song structure while maintaining lyric quality and melody alignment.
Human Evaluation Results
For melody-alignment quality, Reffly achieves the highest scores in prosody, singability, and intelligibility. Lyra performs adequately but falls short compared to Reffly, as it does not align prominent words and notes during generation. SongMASS and GPT-4 have much lower scores, suggesting that their lyrics may not fit well with the melody. This indicates that Reffly excels in generating lyrics that align well with the melody are easy to sing.
For text quality, Reffly scores the highest in both creativity and coherence, indicating its ability to generate lyrics that are both creative and contextually consistent. While GPT-4 performs reasonably well in text quality, its musicality remains poor. The other models score low in coherence, suggesting their lyrics may lack logical progression and contextual consistency.
5.3 Result of Structure-Aware Generation
Among all baselines, only GPT-4 has the capability of generating full-length structured lyrics. Table 4 present the result of structure-aware generation task. Reffly demonstrated a 44% improvement on structural clarity. Figure  10 shows the generated lyrics’s clear verse-chorus-verse-chorus structure.
5.4 Result of Song Translation
Similar to the previous task, only GPT-4 has the capability of song translation, so it is our only baseline. Table 3 presents the evaluation result. Reffly demonstrates a remarkable on average 23% increase in all human evaluation metrics. This results suggest Reffly significantly enhances the quality, coherence, and melody-alignment of generated lyrics compared to GPT-4, making it more suitable for practical applications in song translation.
Interestingly, although GPT-4 generally has stronger translation abilities compared to LLaMA2-13b, Reffly outperforms GPT-4 in translation quality by 14%. This suggests that successful translation in the lyrics-writing context requires not only high text quality, but also an emphasis on how well the lyrics sound when singing, as shown in Figure 4
Figure 4: The output of different models given the same input draft. Reffly is the only model that aligns lyrics with the melody while preserving the original meaning. Other models produce unsingable or low-prosody lyrics (introduced in  Section 2.1). The orange box highlights the lyrics that is not singable. For example, in SongMASS generated lyrics, ‘a-round’, a two-syllable word, is mapped to one musical note, making it hard to sing. The purple box highlights important words that failed to map to a prominent musical note (low prosody), which would disrupt the expressiveness of lyrics. Listen to the audios in the demo page for an intuitive sense.
6 Case Study
We conducted a case study to better understand the advantages of Reffly compared to baselines. An exemplary generated output is shown in Figure 4.
Musicality: Our model generates melody-aligned lyrics while preserving the original draft’s meaning. Unlike the baselines, which overlook the importance of mapping prominent words to prominent notes, our approach ensures that the melody emphasizes these words. In Figure 4, the purple boxes highlight important words that are failed to map with prominent notes, and only Reffly generate melody-aligned lyrics. In addition, SongMASS and ChatGPT-4 generate unsingable lyrics, indicated by yellow box.
Revision capability: Reffly rephrases sentence structures or modifies words, adding ties to ensure prominent word-note alignment. For example, in Figure 4, it rephrased ‘Your calm face reflecting vibrant colors’ into ‘eyes of peace, a canvas of hues’, and ‘You can follow my steps’ into ‘You can follow my footsteps". In both cases, the original meaning is retained. Furthermore, Lyra and SongMASS could not generate coherent lyrics. Although both Reffly and Lyra generate lyrics line-by-line, Reffly produces coherent lyrics due to our training strategy that considers the context of previously generated lines.
6.1 Ablation Study
We conduct ablation study to validate each component in Reffly. A qualitative example in Figure 7 compares different model variations. Reffly-S omits the candidate selection algorithm, and Reffly-I excludes the instruction component in training. Table 1 indicates Reffly-S perform better than Reffly-I, but is prosody score is lower comparing to Reffly. Removing both candidate selection and instruction mechanism would cause a noticeable performance drop.
7 Related work
Controllable lyrics generation
While various M2L generators have been developed to follow control signals like music scores or themes, none provide full control over the generated content. Fan et al. (2019) employed control mechanisms to generate lyrics based on specific topics, Tian et al. (2023) used keywords and genre to control the content, and Saeed et al. (2019), used music audio to condition the generation process. Other approaches have incorporated stylistic elements, such as rhyme schemes and meter, to influence the lyrical output (Potash et al., 2015; Zhang et al., 2022c; Qian et al., 2023). Despite these efforts, the control is typically limited to broad thematic or stylistic aspects, and lacks fine-grained control such as sentence-level semantics and note-level music alignment with the score. To the best of our knowledge, Reffly is the first to provide such capabilities.
Melody-Lyrics alignment
LLMs have proven effective in the M2L generation task, with various attempts to integrate music representation (Zhang et al., 2022c; Lee et al., 2019; Qian et al., 2022). For example, Sheng et al. (2021) applied two transformers for cross-attention between lyrics and melody; Tian et al. (2023) considered duration of musical note in decoding stage, aligning stressed syllables with longer musical notes. However, existing models fail to align prominent words with prominent notes, resulting in poor prosody. We propose a novel heuristic for melody-lyrics alignment, achieving a 26% improvement. Reffly leverages this approach to enhance lyric generation and emotional expression.
8 Conclusion
In this work, we introduced Reffly, the first revision framework to generate high-quality song lyrics from plain text drafts while retaining their original meaning. Furthermore, to enhance the lyrics-melody alignment, we designed a heuristic to identify prominent notes in the melody and important words in lyrics. Using these features and song structure information, we developed an instruction-based mechanism to guide LLMs in generating singable lyrics with key words mapped to prominent notes. Finally, we show the diverse applicability of Reffly demonstrates diverse applicability: it generates full-fledged lyrics with specific structures from scattered user input and is the first to facilitate lyrics translation.
Limitation
The limitation of our work include: 1) In this work, we use a rule-based method to identify important words in lyrics, specifically, nouns, verbs, and adjectives. Future work could investigate more nuanced definitions of important words. 2) Similarly, Our method for extracting prominent notes considers only two levels: prominent notes and other notes. While this simple approach has yielded satisfying results, exploring more fine-grained categories could potentially enhance performance. 3) Although our approach preserves the original meaning of the lyrics, the genre of the resulting lyrics largely depends on the training dataset. Future research could aim to provide more control over the genre.
Appendix A Exemplary Data
A.1 Example of full-length Song
Figure 5: Example song with verse-chorus-verse-chorus structure
A.2 Example of validation dataset
Figure 6: Exemplary data points in validation datasets, where experts annotate the ground truth prominent notes
Figure 7: Example song
Appendix B Human Evaluation Details
Human annotators are paid with \$ 20 per hours.Note that the use of an external voice synthesizer with limited quality impacts the average score of the human evaluation for the Arbitrary Generation task. Despite this, our method outperforms all baselines, with an inter-annotator agreement of 0.69, further demonstrating the effectiveness of our Reffly. Given the current limitations of open-source AI singing voice generation models, the quality of the singing in the generated songs may not meet human standards, which can lead to lower scores. It is important to note that, under the same test settings, our model performs significantly better than all the baselines in every task setting.
B.1 Qualification Task
To evaluate the Turkers’ expertise in the field, we designed a task that included the initial verse from 9 different songs, each with ground-truth labels. These songs were chosen with care to avoid unclear cases, allowing for a clear assessment of quality. The selected songs were those whose scores showed a strong correlation with the ground-truth labels. We select 49 qualified annotators out of 87 annotators, based on Pearson correlation metric. The average inter-rater agreement in therms of Pearson correlation among qualified annotators was 0.43.
B.2 Annotation Task
We present the original survey, including evaluation instructions and the annotation task, in Figure 11 through Figure 14. Figure 11, Figure 12, and Figure 13 outline task instructions, defining each metric—intelligibility, singability, prosody, coherence, creativity—and accompanied by examples of good and bad lyrics for each criterion. Figure 14 display the actual annotation task.
Appendix C Experiments Details
C.1 Details regarding to baseline construction
ChatGPT
We used ChatGPT-4-turbo as the base model to construct this baseline. In order to make this baseline to be fair, we tried our best to prompt ChatGPT-4. Firstly, we use 2-shot manner to prompt it: we provide two golden exemplary revision example every time. To make sure that ChatGPT have the same information that Reffly has, we provided lyrics and the corresponding serialized score using music21, a format that zero-shot ChatGPT could understand. This score encompasses every detail about the music, including rhythm, pitch, and time signature. Note that extracting additional details, such as the position of prominent notes, would require the prominent note extractor from Reffly’s framework. Our objective is to use ChatGPT-4 as a baseline, not to replace Llama2-13b as a revision module.
Lyra
Since the original Lyra paper (Tian et al., 2023) used GPT-2 as the base model, in order to make the comparison fair, we re-implemented the Lyra using Llama-2-13b. When doing the experiments, we use the exact same lyrics drafts as Reffly, which is generated from a collected user prompt. Since Lyra requires keywords as inputs to generate each sentence, we use Yake (Campos et al., 2020) to extract three keywords from the lyrics draft, as the same setting as the original paper.
C.2 Example of the interface used to collect scattered user input
The Figure 8 illustrates the interface of our input-to-draft model. Initially, the user’s requirements are extracted from the prompt using few-shot LLaMA2-13b with intent extraction examples. The extracted requirements are then presented to the user for confirmation, after which a draft is generated based on the confirmed requirements using LLaMA2-13b.
Note that we use the same revision model in all of arbitrary generation, full-length generation, and song translation. Only the input lyrics draft is different, which are generated by LLaMA2-13b in few-shot manner.
Figure 8: Interface used for input to draft
C.3 Computation cost
As illustrated in Figure 2, we “restart” if the all of the outcome of Step 3.1 is unsigable. Restart happens when the length of lyrics draft is too different from the length of music constraint. When rephrasing, we rephrase the original draft so that the length of draft is closer to the length of music constraint. When doing the experiment, all of the lyrics are generated within 3 iterations.
Appendix D Reffly Details
D.1 candidate selection algorithm
The candidate selection algorithm refines a list of lyric candidates to select the best match based on melody constraints, calculating ties and similarities to ensure a singable output.
Figure 9: Training data example
D.2 Training Data Construction
Construct Input Data
We use ChatGPT 3.5 to rephrase sentences either trivially or non-trivially (50% vs. 50%). Trivial rephrasing changes only a few words without altering the sentence structure. The rephrased lyric is then used as the input for revision module training. Additionally, we provide instructions to better guide LLMs revise lyric that better align with melody. The instructions contain information about 1) the phoneme of each word sourced from the CMU pronunciation dictionary, 2) the number of syllables in both the original and rephrased sentences, and 3) guidelines for modifying the lyric drafts, such as making the sentence shorter or longer.
Generate Pseudo Music Constraint
The final piece of input data is music constraint. Due to the lack of melody-aligned data, we generate a pseudo music constraint for lyric dataset, independent of its melody. This approach addresses both the scarcity and copyright issues associated with aligned data. We assume that the lyrics in the training data exhibit good prosody and are singable. we assign a special token, symbolizing a pseudo note, to each syllable in the sentence (singability assumption). If the syllable associated with the special token occupies the stress position of a prominent word, the token denotes an prominent note, otherwise it represents a less prominent note (good-prosody assumption).
Assemble Output Data
The original lyric, prior to rephrasing in the first step, is used as the output for revision module training. Additionally, we craft an "explanation" paragraph help LLMs revise the lyric by breaking down the revision task into multiple simpler sub-tasks ( more detail see Figure 3).
After processing, the inputs consist of the rephrased sentence, the song title, pseudo music constraints, the original lyrics, the song structure, and specific instructions. The output includes the original lyrics accompanied by an explanation.
Exemplary training data point
Figure 9 shows an example training data point constructed using the pipeline introduced in Figure 3. The original lyric is "When you’re listening late at night." We generate a "pseudo melody constraint" based on this lyric, then use ChatGPT to create a rephrased lyric draft, "When you’re tuning in during the nighttime." The model is trained to generate the correct original lyric using the title, lyric draft, pseudo melody constraint, and instruction as input.
Appendix E Music Theory
E.1 Representation in Melody
The representation for a melody is hierarchical. A melody M consists of a series of musical phrase $M=(p_{0},p_{2},..p_{x})$, where x is the total number of musical phrase; Each musical phrase consists of a series of measures $p_{i}=(m_{1},m_{2},...m_{y}),i\in[0,x]$, where y is the total number of measures in i’th musical phrase. Note that $|p_{j}\cap p_{j+1}|\leq 1$. The intersection equals to 1 when a musical phrase end in the middle of a certain measure, so the next musical phrase starts from the same measure. Each measure consists of a series of notes and a corresponding time signature. $m_{k}=(n_{1},n_{2},...,n_{z})$, where z is the total number of notes in measure $m_{k}$. Each note has four component: pitch, duration, offset, and tie. Pitch represents the highness/lowness of a note; duration is the length of the note; offset is the beat when this note starts in its measure; and tie can be start (a tie starts from this note), or continue (in between of a tie), or end (a tie ends at this note).
E.2 Prominent note extraction heuristic details
Inspired by prior research in music theory Caroline Palmer (2006), we develop a more comprehensive heuristics to identify prominent musical notes based on three fundamental characteristics of music:
  1. 1.
    Time Signature: This characteristic provides a structured framework that dictates how beats are grouped and accented within each measure. We identify notes that fall on strong beats or downbeats as prominent notes.
  2. 2.
    Rhythm: For this characteristic, we specifically examine syncopation, a musical technique that shifts emphasis to beats or parts of a beat where it is not usually expected. This technique breaks the conventional rhythmic pattern by highlighting off-beats or weaker beats within a measure. Notes that are accentuated using this technique are identified as prominent notes.
  3. 3.
    Pitch: From this characteristic, we particularly focus on pitch jump. Large pitch jumps contribute to contrast and variety in the melody line, thereby making notes with significant pitch jumps more conspicuous. We classify notes that exhibit significant pitch jumps as prominent notes.
Melody Melody is a sequence of musical tones, consisting of multiple musical phrases that can be further decomposed into timed musical notes. Each musical note has two independent aspect: pitch and duration. Pitch refers to the perceived highness or lowness of a sound; duration refers to the length of time that a musical note is held or sustained.
Time signature time signature organizes the rhythm and provides a framework for how the beats are grouped and accented within each measure. A time signature is represented by two numbers, one stacked on top of the other: the top number indicates the number of beats in each measure; the bottom number indicates the duration value that represents one beat. For example, 4/4 means a quarter note as one beat, 4 beats in a measure. Table 1 shows the stressed location for some commonly-seen time signature. The elements in the list is the number of beat that is stressed. For example, [0,2] means the first and third beats are stressed.
Table 5: Time signatures and their stressed locations
Time Signature Stressed Location
4/4 [0, 2]
3/4 [0]
2/4 [0]
3/8 [0, 2]
6/8 [0, 2]
9/8 [0, 2, 5]
12/8 [0, 2, 5, 8]
Syncopation Syncopation refers to the displacement or shifting of accents or emphasis to unstressed beats. If a note is in unstressed beat with a longer duration than its previous note, then this note, although in unstressed beat, is stressed, or syncopated.
Pitch jump Pitch jump for two consecutive notes is the absolute difference of their pitch value. Larger pitch jumps create contrast and variety within the melody line.
If a note is in a metrical stressed position (indicated by time signature), or it is a syncopation, or it has a pitch jump (larger than average interval), we consider it as a prominent note, otherwise, it is a non-prominent note.
E.3 Prominent words extraction heuristic details
We identify all non-stop words whose part-of-speech are noun, verbs, and adjectives as prominent words.
E.4 Results for heuristics
We provide more details for §$\S$ 5.1 at here. Because our validation dataset only contains ground truth prominent note, we use Yake (Campos et al., 2020) algorithm to extract up to 3 keywords from one lyrics sentence, and treat the extracted keywords that correspond to ground truth prominent notes as ground truth prominent words.
Figure 10: Exemplary generated song with verse-chorus-verse-chorus structure
Figure 11: Human evaluation survey: task instruction
Figure 12: Human evaluation survey: explanation of different metrics 1
Figure 13: Human evaluation survey: explanation of different metrics 2
Figure 14: Human evaluation survey: the annotation task